OpenStudy (anonymous):

Find the indicated nth partial sum of the arithmetic sequence. 1.9, 4.8, 7.7, 10.6, ...., n=20. Show your work.

OpenStudy (amistre64):

you will need to determine the common difference between terms then

OpenStudy (amistre64):

once you determine the common difference "d", there is a general formula for a partial sum\[S_n=\frac{n(a_1+a_n)}{2}\] \[S_n=\frac{n(a_1+(a_1+d(n-1))}{2}\] \[S_n=\frac{n(2a_1+dn-d)}{2}\] since you know the first term a1, and the nth or 20th element all you really needs is the common difference "d"

OpenStudy (anonymous):

so.. \[S _{n}=551\] ? @amistre64

OpenStudy (amistre64):

maybe ... 4.8 1.9 --- 2.9 = d a20 = 1.9+2.9(19) = 57 20(1.9+57) ---------- = 10(58.9) = 589 2

OpenStudy (anonymous):

woopss. somehow i subtracted wrong finding d. ha. okay, i get it now. thanks.

OpenStudy (anonymous):

@amistre64 can you help me with this one?

OpenStudy (amistre64):

1n - 3n

OpenStudy (amistre64):

\[\sum_{n=1}^{299}n-3\sum_{n=1}^{k} n\] 299/3 = 99 + .... so k=99 \[\sum_{n=1}^{299}n-3\sum_{n=1}^{99} n\]

OpenStudy (amistre64):

essentially \[\sum_{n=1}^{299}n-3\sum_{n=1}^{99} n\] \[\sum_{n=1}^{99}n-3\sum_{n=1}^{99} n+\sum_{n=100}^{299}n\] \[-2\sum_{n=1}^{99}n+\sum_{n=100}^{299}n\] \[\frac{99(-2-198)}{2}+\frac{200(100+299)}{2}\]

OpenStudy (amistre64):

or instead of trying to get fancy :)\[\frac{299(1+299)}{2}-3\frac{99(1+99)}{2}\] \[299(150)-3(99)(50)\]

OpenStudy (anonymous):

so the sum of the integers that are not multiples of 3 is 30,006?

OpenStudy (amistre64):

14850

OpenStudy (anonymous):

I'm lost.

OpenStudy (amistre64):

3(99) = 297 so all the multiples of 3 from 1 to 99 are subtracted 3( 1 + 2 + 3 + 4+...+99) +99+98+97+96+...+ 1 ------------------------- 3(100+100+100+100+...+100) |------- 99 times----------| which is 2 times more than we want 99(100) divided in half is 99(50)

OpenStudy (amistre64):

we are basically wanting to count the set twice in a convenient manner take the sum of the digits from 1 to 299 1 + 2 + 3 + 4 +...+296+297+298+299 , lets add this to 299+298+297+296+...+ 4 + 3 + 2 + 1 itself backwards ----------------------------------------- 300+300+300+300+...+300+300+300+300 thats 300, 299 times. we only need half of that count

OpenStudy (anonymous):

Yeah.. im still completely lost :/

OpenStudy (amistre64):

do you agree that if we add up all the integers from 1 to 299, twice, we get 300 * 299 ?

OpenStudy (amistre64):

look at the setup i did, and see if you can convince yourself that its true :)

OpenStudy (amistre64):

we can try it on smaller sets if need be 1+2+3 3+2+1 ------ 4+4+4 = 3*4 = 12, which is twice what we need 12/2 = 6 1+2+3 = 6

OpenStudy (amistre64):

2 + 4+ 6+ 8+10 10+ 8+ 6+ 4+ 2 ----------------- 12+12+12+12+12 = 5*12 = 60, which is twice what we need 60/2 = 30 2 + 4 + 6 + 8 + 10 = 30

OpenStudy (anonymous):

i think i know where im getting lost. after you say 1+2+3 3+2+1 ------- then you say 4+4+4 but where did you get this?

OpenStudy (amistre64):

add the columns ...

OpenStudy (amistre64):

3+1 = 4 2+2 = 4 3+1 = 4

OpenStudy (amistre64):

each column amounts to the sum of the first and last numbers of the arithmetic sequence

OpenStudy (anonymous):

oh. okay. i see where you're coming at from that now.

OpenStudy (amistre64):

since the each column is the equal to the sum of the first and last values; we only really need to know the first and last values ... times the number of "numbers" there are. 3 * (1+3) = 3*4 = 12 ... 12/2 = 6 5*(2+10) = 5*12 = 60 .... 60/2 = 30 299(1+299) = 299*300 = 89700 ..... 89700/2 = 44850

OpenStudy (amistre64):

subtract from that the multiples of 3, from 1 to 99

OpenStudy (amistre64):

3 * (1+99) * 99 = 3 * 100 * 99 ..... and divide by 2: 14850 44850 -14850 --------- 30000